Integral (1-cosx)cosec^2xdx please simplify and then send

Grade:12

3 Answers

Arun
6 years ago
int (1 - cos(x))*csc^2(x) dx 

u = 1 - cos(x); du = sin(x) dx; dv = csc^2(x) dx; v = -cot(x) 

-cot(x) + cot(x)*cos(x) + int cot(x)*sin(x) dx 

-cot(x) + cot(x)*cos(x) + int cos(x) dx 

cot(x)*cos(x) + sin(x) - cot(x)
NARENDRA GODARA
6 years ago
Put cosce^2x=1/sin^2x.then,Put (1-cosx)=2sin^2(x/2) in nemurator.after that put sin^2x=(2sin (x/2)cos (x/2))^2.By solving we getIntegration of 1/2 (sec^2(x/2)).which is equal to 1/2 (tan(x/2)) +c
Sara
5 years ago
ur question is:
(1-cosx)/cosec^2x
cosec^2x – cosecxcotx
whose individal integration is:
-cotx+cosecx